Rafiq exercised for 3 / 6 of an hour, while Rohit exercised for 3 / 4 of an hour. Who exercised for a longer time?

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Rafiq exercised = 3 / 6 of an hour
Rohit exercised = 3 / 4 of a hour 3 / 6, 3 / 4
Convert these into like fractions 3 / 6 = (3× 2) / (6 × 2)
= 6 / 12
3 / 4 = (3 × 3) / (4 × 3)
= 9 / 12
Clearly, 9 / 12 > 6 / 12
∴3 / 4 > 3 / 6
​Therefore Rohit exercised for a longer time than Rafiq.

​Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Rafiq exercised = 3/6 of 1 hr
Rohit exercised = 3/4 of 1 hr
For comparison, we have to make denominator equal,
3×2/6×2 and 3×3/4×3
6/12 and 9/12
 
Therefore, Rohit exercised for longer time than Rafiq.
 
Hope it helps!
Thank you ✌️